OpenStudy (anonymous):

Solve for m: m - 14 = 4.

OpenStudy (anonymous):

m = 10 m = -10 m = 18 m = 4/14 answers

OpenStudy (anonymous):

add 14 to both sides

OpenStudy (anonymous):

so i would get m=18??

OpenStudy (anonymous):

YES :D

OpenStudy (anonymous):

you can confirm that 18-4=? should be 4 and that is right!

OpenStudy (uri):

m - 14 = 4. m=4+14 m=18